# A Note on Minimax Robustness of Designs Against Correlated or Heteroscedastic Responses

### 18 Dec 2023 v2

@@ 1 Introduction
Experimental designs are typically derived, or chosen, assuming that the responses will be independent and homoscedastic. As well as being simple, this is almost necessary, unless an alternative covariance structure is known. This is frequently a complicating feature of design theory – until a design is constructed and implemented there are no data which can be used to estimate an alternate structure. There is some consolation however if a proposed design, optimal under an assumption of independence and homoscedasticity, is minimax against a broad class of alternative covariance structures. By this we mean that the maximum loss in such a class is minimized by this design, which is thus robust against these alternatives. In this note we establish that any function of these alternate covariance matrices, which possesses a natural monotonicity property, is maximized within such classes at a scalar multiple of the identity matrix. This can be paraphrased by saying that the ‘least favourable’ covariance structure is that of independence and homoscedasticity. These are eventualities for which the proposed design is optimal; hence it is minimax.
@@ 2 Main result
(i) For the Euclidean norm $\left\|{x}\right\|_{V}=({{x}^{T}{x}})^{1/2}$, the matrix norm $\left\|C\right\|_{M}=\left\|C\right\|_{E}$ is the spectral radius, i.e. the root of the maximum eigenvalue of $C^{T}C$. This is the maximum eigenvalue of $C$ if $C$ is a covariance matrix, i.e. is symmetric and positive semidefinite.
(ii) For the sup norm $\left\|{x}\right\|_{V}=\max_{i}\left|x_{i}\right|$, the matrix norm $\left\|C\right\|_{M}=\left\|C\right\|_{\infty}$ is $\max_{i}\sum_{j}\left|c_{ij}\right|$, the maximum absolute row sum.
(iii) For the 1-norm $\left\|{x}\right\|_{V}=\sum_{i}\left|x_{i}\right|$, the matrix norm $\left\|C\right\|_{M}=\left\|C\right\|_{1}$ is $\max_{j}\sum_{i}\left|c_{ij}\right|$, the maximum absolute column sum. For symmetric matrices, $\left\|C\right\|_{1}=\left\|C\right\|_{\infty}$.
@@
P1) $\left\|I\right\|_{M}=1$.
P2) For covariance matrices $C$, $\left\|C\right\|_{M}\geq\left\|C\right\|_{E}$.
Property P1) is immediate from (1). For P2), suppose that $\left\|C\right\|_{M}$ is induced by $\left\|\cdot\right\|_{V}$, and that $\lambda_{0}$ is the maximum eigenvalue of $C$, with eigenvector $x_{0}$ normalized so that $\left\|x_{0}\right\|_{V}=1$. Then
@@
For any positive constant $\eta^{2}$, covariance matrix $C$ and norm $\left\|C\right\|_{M}$, define
@@ 3.1 Experimental design in the linear model
Consider the linear model $y=X\theta+\varepsilon$. Suppose that the random errors $\varepsilon$ have covariance matrix $C\in\mathcal{C}_{M}$. If $C$ is known then the ‘Best Linear Unbiased Estimate’ is $\hat{\theta}_{\text{{blue}}}=$ $\left(X^{T}C^{-1}X\right)^{-1}X^{T}C^{-1}y$ and there is an extensive design literature - see Dette et al. (2015) for a review. In the more common case that the covariances are only vaguely known, or perhaps only suspected, it is more usual to use the ordinary least squares estimate $\hat{\theta}_{\text{{ols}}}$, design as though the errors are uncorrelated, and hope for the best. An implication of the results of this section is that, in a minimax sense, that approach can be quite sensible.
In the classical ‘alphabetic’ design problems, one seeks to minimize a function $\Phi$ of the covariance matrix of the regression estimates. Let $\xi_{0}$ be the minimizing design, under the possibly erroneous assumption of uncorrelated, homoscedastic errors. Assume that under $\xi_{0}$the moment matrix $X^{T}X$ is non-singular. Then the covariance matrix of $\hat{\theta}_{\text{{ols}}}$ is
$\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C)=\left(X^{T}X\right)^{-1}X^{T}CX\left(X^{T}X\right)^{-1}.$ (4)
Suppose that $0\preceq C_{1}\preceq C_{2}$, so that $C_{2}-C_{1}=A^{T}A$ for some $A$. Then
$\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C_{2})-\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C_{1})=B^{T}B\succeq 0,$
for $B$ $=AX\left(X^{T}X\right)^{-1}$, hence $\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C_{1})\preceq\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C_{2}).$ It follows that if $\Phi$ is non-decreasing in the Loewner ordering, then $\mathcal{L}(C)=\Phi\{\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C)\}$ is also non-decreasing and the conclusions of the lemma hold. Then as at Remark 1, $\xi_{0}$ is a minimax design – it minimizes the maximum loss as the covariance structure varies over $\mathcal{C}_{M}$.
@@
(iv) $\Phi\left(\Sigma\right)=tr\left(L\Sigma\right)$ for $L\succeq 0$, corresponding to ‘L-optimality’ and including ‘I-optimality’ – minimizing the integrated variance of the predictions;
thus the designs optimal under any of these criteria are minimax.
@@
As a particular case, assume first that the random errors are homoscedastic but are possibly serially correlated, following an ma(1) model with corr$(\varepsilon_{i},\varepsilon_{j})=\rho I\left(\left|i-j\right|=1\right)$ and with $\left|\rho\right|\leq\rho_{\max}$. Then under this structure $C$ varies over the subclass $\mathcal{C}^{\prime}$ of $\mathcal{C}_{\infty}$ defined by $c_{ij}=0$ if $\left|i-j\right|>1$ and $\left\|\boldsymbol{C}\right\|_{\infty}\leq\sigma^{2}\left(1+2\rho_{\max}\right)$, which we define to be $\eta^{2}$. Since $\eta^{2}I_{N}\in\mathcal{C}^{\prime}$, part (ii) of the lemma applies and yields that $\xi_{0}$ is a minimax design in $\mathcal{C}^{\prime}$ and with respect to any of the alphabetic criteria above. If the errors are instead heteroscedastic, then $\sigma^{2}$ is replaced by the maximum of the assumed variances.
@@
It is known that the eigenvalues of an ar(1) autocorrelation matrix with autocorrelation parameter $\rho$ are bounded, and that the maximum eigenvalue $\lambda_{N}\left(\rho\right)$ has $\lambda_{N}^{\ast}=$ $\max_{\rho}\lambda_{N}\left(\rho\right)>\lambda_{N}\left(0\right)=1$. See for instance Trench (1999, p. 182). Then, again under homoscedasticity, the covariance matrix $C$ has $\left\|C\right\|_{E}\leq\sigma^{2}\lambda_{N}^{\ast}$, and a design optimal when $\rho=0$ is minimax in the subclass of $\mathcal{C}_{E}$ defined by the autocorrelation structure and $\eta^{2}=\sigma^{2}\lambda_{N}^{\ast}$.
@@ 3.2 Designs robust against model misspecifications
$Y\left(x\right)=f^{T}\left(x\right)\theta+\psi\left(x\right)+\varepsilon,$ (5)
@@
$\mathcal{I}\left(\psi,\xi\right)$ $=$ $\sum_{x\in\chi}E[f^{T}\left(x\right)\hat{\theta}-E\{Y\left(x\right)\}]^{2},$ (7a)
$\mathcal{D}\left(\psi,\xi\right)$ $=$ $[\det E\{(\hat{\theta}-\theta)(\hat{\theta}-\theta)^{T}\}]^{1/p},$ (7b)
@@
$\mathcal{I}\left(\psi,\xi,C\right)=\sum_{x\in\chi}f^{T}\left(x\right){\textsc{cov}}(\hat{\theta}\mid C)f\left(x\right)+\sum_{x\in\chi}\left\{f^{T}\left(x\right)d_{\psi}d_{\psi}^{T}f\left(x\right)+\psi^{2}\left(x\right)\right\}.$ (8)
@@
$\max_{\psi,C}\mathcal{I}\left(\psi,\xi,C\right)=\max_{C\in\mathcal{C}_{M}}\sum_{x\in\chi}f^{T}\left(x\right)\text{{cov}}(\hat{\theta}\mid C)f\left(x\right)+\max_{\psi\in\Psi}\sum_{x\in\chi}\left\{f{}^{T}\left(x\right)d_{\psi}d_{\psi}^{T}f\left(x\right)+\psi^{2}\left(x\right)\right\}.$
As in §3.1, and taking $L=\sum_{x\in\chi}f\left(x\right)f^{T}\left(x\right)$ in (iv) of that section, for $C\in\mathcal{C}_{M}$ the first sum is maximized by a multiple $\eta^{2}$ of the identity matrix, and then the remainder of the minimax problem is that which was solved in Wiens (2018). The minimax designs, termed ‘I-robust’ designs, obtained there do not depend on the value of $\eta^{2}$, and so enjoy the extended property of minimizing $\max_{\psi,C}\mathcal{I}\left(\psi,\xi,C\right)$ for $C\in\mathcal{C}_{M}$.
Now consider (7b). The analogue of (8) is $\mathcal{D}\left(\psi,\xi,C\right)=[\det\{\text{{cov}}(\hat{\theta}\mid C)+d_{\psi}d_{\psi}^{T}\}]^{1/p}.$ Since cov$(\hat{\theta}\mid C)+d_{\psi}d_{\psi}^{T}$, hence its determinant, is non-decreasing in the Loewner ordering, $\mathcal{D}\left(\psi,\xi,C\right)$ is maximized, for $C\in\mathcal{C}_{M}$, by a multiple of the identity matrix. The rest of the argument is identical to that in the preceding paragraph, and so the ‘D-robust’ designs obtained in Wiens (2018) also minimize $\max_{\psi,C}\mathcal{D}\left(\psi,\xi,C\right)$ for $C\in\mathcal{C}_{M}$.
@@ 3.3 Designs for nonlinear regression models
In the nonlinear regression model $y=f\left(x;\theta\right)+\varepsilon$, the goal of a design is typically the minimization of some function of the covariance matrix of $\hat{\theta}$, after a linearization of the model. When the errors have covariance $C$ the target of this minimization continues to be given by the right hand side of (4), but with the matrix $X$ replaced by the gradient $F\left(\theta\right)=\{\partial f\left(x_{i};\theta\right)/\partial\theta_{j}\}_{i,j}$, and with $F\left(\theta\right)$ evaluated at an initial parameter estimate, or at a previous estimate in an interative estimation scheme. Then the results of §3.1 continue to hold, and a design optimal for the loss functions given there, under independence and homoscedasticity, is minimax. A caveat, however, is that optimal design theory for nonlinear models is much less well-developed, and more model specific, than that for linear models. See Bates and Watts (1988) and Hamilton and Watts (1985) for discussions.
Acknowledgement

### 02 Jan 2024 v3

@@ 1 Introduction
Experimental designs are typically derived, or chosen, assuming that the responses will be independent and homoscedastic. As well as being simple, this is almost necessary, unless an alternate covariance structure is somehow known. This is frequently a complicating feature of design theory – until a design is constructed and implemented there are no data which can be used to estimate the model. There is some consolation however if a proposed design, optimal under an assumption of independence and homoscedasticity, is minimax against a broad class of alternate structures. By this we mean that the maximum loss in such a class is minimized by this design, which is thus robust against these alternatives. In this note we establish that any function of covariance matrices, possessing a natural monotonicity property, is maximized within such classes at a scalar multiple of the identity matrix. This can be paraphrased by saying that the ‘least favourable’ covariance structure is that of independence and homoscedasticity. These are eventualities for which the proposed design is optimal; hence it is minimax.
@@ 2 Main result
(i) For the Euclidean norm $\left\|{x}\right\|_{V}=({{x}^{{\mathrm{\scriptscriptstyle T}}}{x}})^{1/2}$, the matrix norm is denoted $\left\|C\right\|_{E}$ and is the spectral radius, i.e. the root of the maximum eigenvalue of $C^{{\mathrm{\scriptscriptstyle T}}}C$. This is the maximum eigenvalue of $C$ if $C$ is a covariance matrix, i.e. is symmetric and positive semidefinite.
(ii) For the sup norm $\left\|{x}\right\|_{V}=\max_{i}\left|x_{i}\right|$, the matrix norm $\left\|C\right\|_{\infty}$ is $\max_{i}\sum_{j}\left|c_{ij}\right|$, the maximum absolute row sum.
(iii) For the 1-norm $\left\|{x}\right\|_{V}=\sum_{i}\left|x_{i}\right|$, the matrix norm $\left\|C\right\|_{1}$ is $\max_{j}\sum_{i}\left|c_{ij}\right|$, the maximum absolute column sum. For symmetric matrices, $\left\|C\right\|_{1}=\left\|C\right\|_{\infty}$.
@@
(P1) $\left\|I\right\|_{M}=1$.
(P2) For covariance matrices $C$, $\left\|C\right\|_{M}\geq\left\|C\right\|_{E}$.
Property P1 is immediate from (1). For P2, suppose that $\left\|C\right\|_{M}$ is induced by $\left\|\cdot\right\|_{V}$, and that $\lambda_{0}$ is the maximum eigenvalue of $C$, with eigenvector $x_{0}$ normalized so that $\left\|x_{0}\right\|_{V}=1$. Then
@@
For $\eta^{2}>0$, covariance matrix $C$ and induced norm $\left\|C\right\|_{M}$, define
@@ 3.1 Experimental design in the linear model
Consider the linear model $y=X\theta+\varepsilon$. Suppose that the random errors $\varepsilon$ have covariance matrix $C\in\mathcal{C}_{M}$. If $C$ is known then the ‘Best Linear Unbiased Estimate’ is $\hat{\theta}_{\text{{blue}}}=$ $\left(X^{{\mathrm{\scriptscriptstyle T}}}C^{-1}X\right)^{-1}X^{{\mathrm{\scriptscriptstyle T}}}C^{-1}y$ and there is an extensive design literature - see Dette et al. (2015) for a review. In the more common case that the covariances are only vaguely known, or perhaps only suspected, it is more usual to use the ordinary least squares estimate $\hat{\theta}_{\text{{ols}}}$, design as though the errors are uncorrelated, and hope for the best. An implication of the results of this section is that, in a minimax sense, that approach can be sensible.
In the classical ‘alphabetic’ design problems, one seeks to minimize a function $\Phi$ of the covariance matrix of the regression estimates. Let $\xi_{0}$ be the minimizing design, under the possibly erroneous assumption of uncorrelated, homoscedastic errors. Assume that under $\xi_{0}$the moment matrix $X^{{\mathrm{\scriptscriptstyle T}}}X$ is non-singular. Then the covariance matrix of $\hat{\theta}_{\text{{ols}}}$ is
$\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C)=\left(X^{{\mathrm{\scriptscriptstyle T}}}X\right)^{-1}X^{{\mathrm{\scriptscriptstyle T}}}CX\left(X^{{\mathrm{\scriptscriptstyle T}}}X\right)^{-1}.$ (4)
Suppose that $0\preceq C_{1}\preceq C_{2}$, so that $C_{2}-C_{1}=A^{{\mathrm{\scriptscriptstyle T}}}A$ for some $A$. Then
$\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C_{2})-\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C_{1})=B^{{\mathrm{\scriptscriptstyle T}}}B\succeq 0,$
for $B$ $=AX\left(X^{{\mathrm{\scriptscriptstyle T}}}X\right)^{-1}$, hence $\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C_{1})\preceq\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C_{2}).$ It follows that if $\Phi$ is non-decreasing in the Loewner ordering, then $\mathcal{L}(C)=\Phi\{\textsc{cov}(\hat{\theta}_{\textsc{ols}}\mid C)\}$ is also non-decreasing and the conclusions of the lemma hold. Then as at Remark 1, $\xi_{0}$ is a minimax design – it minimizes the maximum loss as the covariance structure varies over $\mathcal{C}_{M}$.
@@
(iv) $\Phi\left(\Sigma\right)=tr\left(L\Sigma\right)$ for $L\succeq 0$, corresponding to ‘L-optimality’ and including ‘I-optimality’ – minimizing the integrated variance of the predictions. Thus the designs optimal under any of these criteria are minimax.
@@
As a particular case, assume first that the random errors are homoscedastic but are possibly serially correlated, following an ma(1) model with corr$(\varepsilon_{i},\varepsilon_{j})=\rho I\left(\left|i-j\right|=1\right)$ and with $\left|\rho\right|\leq\rho_{\max}$. Then under this structure $C$ varies over the subclass $\mathcal{C}^{\prime}$ of $\mathcal{C}_{\infty}$ defined by $c_{ij}=0$ if $\left|i-j\right|>1$ and $\left\|{C}\right\|_{\infty}\leq\sigma^{2}\left(1+2\rho_{\max}\right)$, which we define to be $\eta^{2}$. Since $\eta^{2}I_{N}\in\mathcal{C}^{\prime}$, part (ii) of the lemma applies and yields that $\xi_{0}$ is a minimax design in $\mathcal{C}^{\prime}$ and with respect to any of the alphabetic criteria above. If the errors are instead heteroscedastic, then $\sigma^{2}$ is replaced by the maximum of the variances.
@@
It is known that the eigenvalues of an ar(1) autocorrelation matrix with autocorrelation parameter $\rho$ are bounded, and that the maximum eigenvalue $\lambda\left(\rho\right)$ has $\lambda^{\ast}=$ $\max_{\rho}\lambda\left(\rho\right)>\lambda\left(0\right)=1$. See for instance Trench (1999, p. 182). Then, again under homoscedasticity, the covariance matrix $C$ has $\left\|C\right\|_{E}\leq\sigma^{2}\lambda^{\ast}$, and a design optimal when $\rho=0$ is minimax in the subclass of $\mathcal{C}_{E}$ defined by the autocorrelation structure and $\eta^{2}=\sigma^{2}\lambda^{\ast}$.
@@ 3.2 Designs robust against model misspecifications
$Y\left(x\right)=f^{{\mathrm{\scriptscriptstyle T}}}\left(x\right)\theta+\psi\left(x\right)+\varepsilon,$ (5)
@@
$\mathcal{I}\left(\psi,\xi\right)$ $=$ $\sum_{x\in\chi}E[f^{{\mathrm{\scriptscriptstyle T}}}\left(x\right)\hat{\theta}-E\{Y\left(x\right)\}]^{2},$ (7a)
$\mathcal{D}\left(\psi,\xi\right)$ $=$ $[\det E\{(\hat{\theta}-\theta)(\hat{\theta}-\theta)^{{\mathrm{\scriptscriptstyle T}}}\}]^{1/p},$ (7b)
@@
$\mathcal{I}\left(\psi,\xi,C\right)=\sum_{x\in\chi}f^{{\mathrm{\scriptscriptstyle T}}}\left(x\right){\textsc{cov}}(\hat{\theta}\mid C)f\left(x\right)+\sum_{x\in\chi}\left\{f^{{\mathrm{\scriptscriptstyle T}}}\left(x\right)d_{\psi}d_{\psi}^{{\mathrm{\scriptscriptstyle T}}}f\left(x\right)+\psi^{2}\left(x\right)\right\}.$ (8)
@@
$\max_{\psi,C}\mathcal{I}\left(\psi,\xi,C\right)=\max_{C\in\mathcal{C}_{M}}\sum_{x\in\chi}f^{{\mathrm{\scriptscriptstyle T}}}\left(x\right)\text{{cov}}(\hat{\theta}\mid C)f\left(x\right)+\max_{\psi\in\Psi}\sum_{x\in\chi}\left\{f^{{\mathrm{\scriptscriptstyle T}}}\left(x\right)d_{\psi}d_{\psi}^{{\mathrm{\scriptscriptstyle T}}}f\left(x\right)+\psi^{2}\left(x\right)\right\}.$
As in §3.1, and taking $L=\sum_{x\in\chi}f\left(x\right)f^{{\mathrm{\scriptscriptstyle T}}}\left(x\right)$ in (iv) of that section, for $C\in\mathcal{C}_{M}$ the first sum is maximized by a multiple $\eta^{2}$ of the identity matrix, and then the remainder of the minimax problem is that which was solved in Wiens (2018). The minimax designs, termed ‘I-robust’ designs, obtained there do not depend on the value of $\eta^{2}$, and so enjoy the extended property of minimizing $\max_{\psi,C}\mathcal{I}\left(\psi,\xi,C\right)$ for $C\in\mathcal{C}_{M}$.
Now consider (7b). The analogue of (8) is
$\mathcal{D}\left(\psi,\xi,C\right)=[\det\{\text{{cov}}(\hat{\theta}\mid C)+d_{\psi}d_{\psi}^{{\mathrm{\scriptscriptstyle T}}}\}]^{1/p}.$
Since cov$(\hat{\theta}\mid C)+d_{\psi}d_{\psi}^{{\mathrm{\scriptscriptstyle T}}}$, hence its determinant, is non-decreasing in the Loewner ordering, $\mathcal{D}\left(\psi,\xi,C\right)$ is maximized, for $C\in\mathcal{C}_{M}$, by a multiple of the identity matrix. The rest of the argument is identical to that in the preceding paragraph, and so the ‘D-robust’ designs obtained in Wiens (2018) also minimize $\max_{\psi,C}\mathcal{D}\left(\psi,\xi,C\right)$ for $C\in\mathcal{C}_{M}$.
@@ 3.3 Designs for nonlinear regression models
In the nonlinear regression model $y=f\left(x;\theta\right)+\varepsilon$, the goal of a design is typically the minimization of some function of the covariance matrix of $\hat{\theta}$, after a linearization of the model. When the errors have covariance $C$ the target of this minimization continues to be given by the right hand side of (4), but with the matrix $X$ replaced by the gradient $F\left(\theta\right)=\{\partial f\left(x_{i};\theta\right)/\partial\theta_{j}\}_{i,j}$, and with $F\left(\theta\right)$ evaluated at an initial parameter estimate, or at a previous estimate in an iterative estimation scheme. Then the results of §3.1 continue to hold, and a design optimal for the loss functions given there, under independence and homoscedasticity, is minimax. A caveat, however, is that optimal design theory for nonlinear models is much less well-developed, and more model specific, than that for linear models. See Bates and Watts (1988) and Hamilton and Watts (1985) for discussions.
Acknowledgements